OpenStudy (anonymous):

standard form for (3,3) (-6,-6)

Elsa213 (elsa213):

\(\frak\Huge\rlap{\color{Magenta}{^*\star^\heartsuit~Welcome~To~Openstudy}}{\color{Orchid}{\:~~~~~~~~~~Welcome~To~Openstudy^*\star^\heartsuit}}\)\(\hspace{-270 pt}\Huge\color{red}{^*:}\color{gold}{_・゚}\color{lime}{^✧}\color{cyan}{_*}\color{blue}{:}\color{purple}{^・゚}\color{blueviolet}{_✧}\color{magenta}{^*}\color{orchid}{:}\color{red}{_・゚}\color{gold}{^✧}\color{lime}{_*}\color{cyan}{:}\color{blue}{^・゚}\color{purple}{✧}\)

jigglypuff314 (jigglypuff314):

Hello and Welcome to OpenStudy! :) first you would have to make an equation out of those two points do you know how to do that?

jigglypuff314 (jigglypuff314):

first you would find the slope between two points \[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} } \]when given points (x1, y1) and (x2, y2) then get an equation through either point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1) or slope-intercept form: y = mx + b when m = slope and b = y-intercept then from there you can get it into standard form Ax + By = C where A, B, and C are whole numbers by getting the terms with variables all to one side